OpenStudy (anonymous):

Can someone evaluate this? Integrate this function from [0,a] where f(x)dx/f(x)+f(a-x)

OpenStudy (anonymous):

\[\int\limits_{0}^{a}\frac{ f(x)dx}{ f(x)+f(a-x) }\]

OpenStudy (anonymous):

-> Provided that the function is both Continuous and differentiable

OpenStudy (photon336):

@Kainui

OpenStudy (freckles):

try to sub u=a-x

OpenStudy (freckles):

if you don't know where to go from there, show me what you have after doing that.

OpenStudy (freckles):

I will give one more hint just in case you aren't there and when you back I'm not here... \[\text{ If } \int_0^a g(x) dx=\int_0^a h(x) dx \text{ then } \\ 2 \int_0^a g(x) dx=\int_0^a g(x) dx+\int_0^a h(x) dx=\int_0^a (g(x)+h(x) )dx\]

OpenStudy (freckles):

\[\text{ If } 2 \int\limits_0^a g(x) =\int\limits_0^a (g(x)+h(x)) dx \\ \text{ then } \int\limits_0^a g(x) dx= \frac{1}{2} \int\limits_0^a (g(x)+h(x)) dx\]

OpenStudy (freckles):

anyways holler if you get stuck somewhere but please show me your work so I can know where you are stuck like even the substitution thing I asked about earlier